Sol Summary – October 30th


Summary Title: First Full Working Day on Mars
Author’s name: Guy Murphy
Mission Status: Full operational status achieved.
Sol Activity Summary: The morning was spent preparing for the afternoon’s EVA, maintaining the campus tunnels, and on other project and administrative work. During the afternoon the first full science EVA was undertaken with 3 crew members using the rovers. 2 cooked meals were prepared today. A few additional food supplies arrived today.
Look Ahead Plan: The EVA program will continue for the remainder of the week subject to favourable weather conditions.
Anomalies in work: None
Weather: Extremely cold night, cold temperatures but clear skies during the day. A colder night anticipated. No precipitation.
Crew Physical Status: Generally good. One crew member experienced a small cut on one hand from a protruding screw head when opening the RAM hatch. The cut has been cleaned and covered with a band aid. Crew member had already been vaccinated for tetanus.
EVA: First full EVA successfully undertaken.
Reports to be filed: Sol Summary, EVA Report, Journalists Report, Science Report, Operations Reports, Photos.
Support Requested: None. An additional food delivery is expected tomorrow.

Science Report – October 30th

Crew 214 Science Report 30 October 2019
SOL: 3
Crew 214 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Operations: The first of the scientific EVAs was scheduled for the afternoon work period. This involved a 2 hour traverse in the rovers to Tank Wash and south of the Galileo and Cactus Road junctions. At both these locations, relatively flat relatively undisturbed areas were identified and GPS surveyed and photographed so that a 10 square metre grid could be deployed from which samples targeting mircometeorites would be collected.
In preparation for the sampling, during the morning work period, the Science Dome’s large open area was utilized to construct the 1m x 1m grid square boundaries in readiness for deployment.
Concurrent with the preparations for micrometeorite samplings and EVA, tabulation of the baseline food stock macro and micronutrients was undertaken for the waste study.
Closed Loop Food Waste Study: The day’s waste for Sol 3 is: 89 grams; consisting of 293 calories, 1g fat, >1g saturated fat, 555mg sodium, 65g carbohydrates, 6g fibre, 31g sugar, 10g protein.
Additional Activities: Reinforcing of the coverings for the tunnels was undertaken throughout the day as time became available.

Operations Report – October 30th

Crew 214 Operations Report 30-10-2019
SOL: 3
Name of person filing report: David Mateus
Non-nominal systems: Opportunity, suit 10
Notes on non-nominal systems: Oppy brakes do not work and Suit 10 is not charging
Generator: run
Hours run: 14:00
From what time last night: 17:30
To what time this morning: 7:30
List any additional daytime hours when the generator was run: none
Solar— SOC % (Before generator is run at night): 80%
Diesel Reading –73 %
Station Propane Reading – 86 %
Ethanol Free Gasoline – N/A
Water (loft tank): 40 gallons
Water Meter: 1454844 units
Water (static tank) – 301 gallons
Static to Loft Pump used – Yes
Water in GreenHab – 150 gallons
Water in ScienceDome: 0 gallons
Toilet tank emptied: no
Deimos rover used: It is not in the hab
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: No
Sojourner rover used: ASSIGNED TO DIRECTOR
Hours: –
Beginning charge: –
Ending charge: –
Currently charging: –
Spirit rover used: Yes
Hours: (before EVA): 106.1 h
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging):47%
Currently charging: Yes
Opportunity rover used: No
Hours: 65.5 h
Beginning charge: 100%
Ending charge: 100%
Currently charging: Yes
Curiosity rover used: Yes
Hours: 112
Beginning charge: (Before EVA): 100%
Ending charge: (On return from EVA, before recharging):59%
Currently charging: Yes
Notes on rovers: No
ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3) No
Reason for use: None
Oil Added? No
ATV Fuel Used: 0 Gals
# Hours the ATVs were Used today: 0
Notes on ATVs: Assigned to the director
HabCar used and why, where? Used by assistant director to go to town
CrewCar used and why, where? No used
General notes and comments: None
Summary of internet: functional
Summary of suits and radios: used for EVA, Suit 10 is not charging and we need batteries for 6 radios
Summary of Hab operations: All systems are nominal
Summary of GreenHab operations: No used
Summary of ScienceDome operations: All systems are nominal, used for experiment preparations
Summary of RAM operations: No used
Summary of any observatory issues: No used
Summary of health and safety issues: All the crew are healthy
Questions, concerns and requests to Mission Support: We need new batteries for six radios.

Journalist Report – October 30th

Wed 30 oct Sol 3
After a particularly cold night we wake to a clear sunny day. This is in some ways our first ‘normal’ day, we have finished bedding in and training up, now for normal life, on Mars.
We have recieved CapCom permission to have an EVA in the afternoon, so we will spend the morning preparing for that. In the meantime, good news, food arrived from Hanksville, however not including the milk and pasta we were hanging out for.
So the morning was spent tying loops on bits of string for use out in the field, as a grid reference system for the micrometeroite experiment. This consists of a 10 metre by 10 metre grid in 1 metre steps. A magnet is used to attract meteorites, if any, from the ground, and the total from each square metre is bagged for offline study. Sounds simple, but executing all these steps in suits can be quite tricky.
In the morning others were busy tying down the tarpaulins sheltering the ‘tunnels’ that connect the various parts of the MDRS campus. This is necessary due to the occasional high winds in these parts. Maybe not such a problem on Mars since the atmosphere is about 1% the density of ours, the gale in ‘The Martian’ movie notwithstanding.
In the afternoon an expedition of 3 Martians and 2 rovers headed north up the track to suss out which of four potential meteorite sites were best for setting up the grid. Out in the backblocks they came across a cougar print, not exactly to be expected on Mars, but a fascinating finding anyway!

EVA Report – October 30th

Crew 214 EVA Report 30-10-2019
EVA # 3
Author: Andrew Wheeler
Purpose of EVA: Geological Reconnaissance
Start time: 1330
End time: 1540
Narrative: This EVA is the first of the scientific activities. Locations for deployment of a sampling grid necessary for collection of magnetic material from the top 1 cm of the regolith that could be identified as micrometeorites were visited and evaluated. One corner of potential grid locations were GPS surveyed and photographed for later layout of the grid. There are two locations near Tank Wash off Cow Dung Road (1) and two locations southeast of the junction of Galileo Road and Cactus Road (2). They are as follows:
Tank Wash #1 518472E 4253624N (preference #4)
Tank Wash #2 518477E 4253767N (preference #1)
Cactus Road #1 519570E 4251974N (preference #3)
Cactus Road #2 519958E 4251995N (preference #2)
Destination: Drove from the hab along Cow Dung Road to Tank Wash and then returned to Galileo Road and the junction it has with Cactus Road.
Coordinates (use UTM NAD27 CONUS): (1): 518500E 4253800N (2): 520100E 4251500N
Participants: David Mateus, Guy Murphy, Andrew Wheeler
Road(s) and routes per MDRS Map: Cow Dung Road, Galileo Road, Cactus Road

Copyright © The Mars Society. All rights reserved. | Main Site